# October   ||   2024 AI-Driven Control of Bioelectric signalling

### v1

@@ 1.2 Synthetic Biology and Multicellularity
Pezzulo and Levin [Pezzulo2016] advocate for top-down approaches in understanding and controlling complex biological systems, contrasting with traditional bottom-up, molecular-centric frameworks. Their study emphasizes the benefits of higher-level organizational strategies, akin to those in physics and engineering, for regulating complex processes such as pattern formation and regeneration. The methodology involves integrating concepts from information theory, control theory, and computational neuroscience to develop models that focus on system-wide goal states and dynamic regulation rather than precise molecular manipulation. Examples include the use of least-action principles in physics and control-theoretic models in neuroscience to achieve system stability and desired outcomes. In developmental biology, the authors propose that targeting large-scale anatomical goals can guide cellular behavior toward desired outcomes through bioelectric signaling and pattern memory. Studies on phenomena like planaria regeneration, where bioelectric interventions led to permanent anatomical changes, support the feasibility of top-down models. The results suggest that focusing on higher-level system organization can provide more effective control over complex biological processes, paving the way for innovative approaches in regenerative medicine and synthetic bioengineering by harnessing goal-directed, system-wide controls over molecular mechanisms.
@@ 1.3 Questions
Bioelectric signals in cells are, again, electrical potentials generated by the movement of ions (such as sodium, potassium, calcium, and chloride) across the cell membrane through ion channels and pumps. Bioelectric signals are represented by the voltage gradients across cell membranes ($V_{mem}$) and play an important role in developmental biology, governing processes such as tissue formation, wound healing, and organogenesis. Recent work by Michael Levin and others has demonstrated that it is possible to manipulate these bioelectric signals, showing the capacity to direct tissue regeneration in organisms like planaria and frogs [Levin2014wh]. Yet, many questions remain unanswered, such as:
@@ 1.4 Research Objectives
For Objective 2, for example, a study in replicability can be conducted following [Pietak2016] where an open-source framework for simulating 2D computational multiphysics problems such as electrodiffusion, electro-osmosis, galvanotaxis, voltage-gated ion channels, gene regulatory networks, and biochemical reaction networks (e.g., metabolism) has been made available through BETSE (BioElectric Tissue Simulation Engine) Github.
@@ 2.1 Current Understanding of Bioelectric Signals
Bioelectric signals are fundamental to the regulation of cellular and developmental processes. They are generated by the movement of ions across cellular membranes, creating voltage gradients ($V_{\text{mem}}$) that influence cell behaviour. Recent research has begun to map the bioelectric patterns associated with various cellular states and developmental processes, revealing that these patterns play a crucial role in tissue formation, regeneration, and organogenesis [Levin2009, Levin2014wh, adams2013endogenous].
Just as the genome has been sequenced to understand genetic contributions to biology, efforts have been made to map bioelectric patterns in organisms. Techniques such as voltage-sensitive dyes and fluorescent reporters have allowed visualization of membrane potential distributions in living tissues. Studies have shown that specific bioelectric patterns correlate with particular developmental outcomes. For example, the formation of the anterior-posterior axis in Xenopus laevis embryos is regulated by bioelectric cues [adams2013endogenous, Pai2012].
However, unlike the genome, the bioelectric "map" is dynamic and context-dependent. The spatial and temporal variability of bioelectric signals across different cell types and developmental stages makes comprehensive mapping challenging. Current maps are incomplete and often organism-specific, limiting the generalization of findings [McCaig2005, Sundelacruz2009].
Cells interpret bioelectric signals through voltage-gated ion channels, transporters, and voltage-sensitive signalling pathways. Changes in membrane potential can influence gene expression, cell proliferation, differentiation, migration, and apoptosis. For instance, depolarization of membrane potential has been shown to promote stem cell proliferation, while hyperpolarization can induce differentiation [Levin2009, Levin2014wh, Blackiston2011].
Despite these insights, the precise mechanisms by which cells transduce bioelectric signals into specific biochemical responses remain poorly understood. The complexity arises from the interplay between electrical, chemical, and mechanical signals within the cellular microenvironment. Additionally, the same bioelectric signal can elicit different responses depending on the cell type and its developmental context [Levin2014wh].
@@ 2.2 Bioelectricity and its Role in Topological Control
Bioelectricity, the result of ion flow across cellular membranes, acts as a fundamental signal that coordinates cell behaviour, gene expression, and tissue organization. These voltage gradients serve as a communication medium that enables cells to “know” their relative position within an organism, guiding processes like cell proliferation, differentiation, and morphogenesis [Levin2014wh, Whited2019]. Levin’s experiments with planaria have demonstrated that by manipulating these bioelectric patterns, researchers can direct tissue outcomes - regenerating a planaria’s head or a tail based solely on the bioelectric state [Beane2011-yb]. These findings suggest that bioelectric signals function as a form of "code" that governs tissue identity, growth, and maintenance.
Additionally, bioelectric signals have been shown to play a role in broader topological manipulations. For instance, in Xenopus laevis (the African clawed frog), bioelectric manipulations have been used to induce the formation of eyes in unconventional locations, illustrating the potential of bioelectric control to not only regenerate tissues but also alter and transform them in novel ways [Levin2007-eb]. This points to the possibility that the bioelectric "fingerprint" behind an organ or tissue could be decoded, allowing for manipulation of $V_{mem}$ at a fundamental level, potentially enabling the full recovery of damaged organs, such as in cardiac arrest victims, complete overhaul of malfunctioning tissues like cancer, and perhaps even the transformation of tissues into entirely novel structures [Davies2020].
Based on the work of Souidi et al., we take the drosophila fly as our goal system for modelling heart disease and recovery induced by the proposed DRL system [Souidi2021].
@@ 2.3 Biomedical and Engineering Implications
The biomedical implications of cracking the bioelectric code are many. By understanding and controlling the high-level triggers of morphogenetic subroutines, it might seem like a giant leap in imagination to picture a moment in time when we’ll be able to induce complex regenerative outcomes such as limb and organ regeneration, repair of birth defects, and even the prevention and reprogramming of diseases like cancer via bioelectric pathways. But these processes have been made possible to some extent already in vivo, both using human cells as well as in complex organisms, such as frogs [Levin2014wh].
Bioelectric manipulation has shown promise in modulating the tumor environment, potentially suppressing or redirecting cancerous cell behaviour [Chernet2013], while optogenetic control of voltage gradients has successfully triggered regeneration in non-regenerative species [Whited2019]. Simulations, like those done with BETSE, allow us to predict and manipulate bioelectric patterns across tissues, aiding in the exploration and improvement over the control of cellular behaviour during morphogenesis [Pietak2016]. Moreover, advances in synthetic biology have begun to explore self-assembling tissues with designed bioelectric circuits, paving the way for the engineering of new tissues and organs, as well as the development of novel therapeutic strategies for regenerative medicine, such as in victims of frostbite, and synthetic life [Davies2020].
@@ 3.2 Reinforcement Learning Framework
DRL has achieved remarkable success in mastering complex games with very high dimensional state representations, as demonstrated by algorithms like AlphaGo [Silver2016], AlphaZero [silver2017mastering], MuZero [Schrittwieser_2020], OpenAI Five for Dota 2 [openai2019dota], and AlphaStar for StarCraft II [vinyals2019grandmaster]. These algorithms have showcased the ability of RL agents to learn optimal strategies in environments with vast state and action spaces, long time horizons, and partial observability [deCarvalho2024deep] (See Appendix LABEL:sec:drl for a deep dive into the algorithms).
@@ 3.2.3 Modelling Heart Topology and Bioelectric Patterns
In bioelectric experiments conducted by Michael Levin’s team, similar principles are applied to study pattern formation in regenerating tissues, such as frog embryos. Levin’s group uses voltage-sensitive dyes, such as DiBAC4(3) and CC2-DMPE, to detect membrane potential differences across cells. These dyes change their fluorescence based on the membrane potential, providing real-time spatial and temporal mapping of the bioelectric state in both normal and abnormal tissues. Through fluorescent microscopy, these bioelectric states are visualized and quantified, allowing the identification of deviations in abnormal tissues compared to healthy ones [adams2013endogenous].
To induce a normal bioelectdrcfx ric pattern in an abnormal frog embryo, ion channels are pharmacologically modulated or optogenetics is used. Optogenetics involves light-sensitive ion channels introduced into the tissue, normally through CRISP-R, enabling precise control over membrane potential with light. By manipulating the ion flow and thus the membrane potential, researchers have been able to correct the bioelectric patterns in the abnormal tissue to match those of the healthy, normal tissue. For instance, an abnormal head pattern in a frog can be corrected by restoring the normal bioelectric gradient, guiding the cells to form a correctly structured and functional head [adams2013endogenous].
This approach, known as bioelectric modulation, effectively uses the bioelectric pattern as a template for normal tissue development and regeneration. It demonstrates that bioelectric states serve as key regulators of tissue and organ formation, enabling control over developmental processes through targeted modulation of membrane potentials [adams2013endogenous].
@@ 4 Causal Inference in Deep Reinforcement Learning Framework
Understanding the causal relationships between bioelectric signals and cellular behaviours could aid control and manipulation of morphogenetic processes. Integrating Judea Pearl’s do-calculus into our DRL framework could allow for more precise causal inference, helping with the agent’s decision-making processes [pearl_book, Pearl2010, Pearl1995]. This graphical representation can also function as an interpretability and explainability tool, regardless of its impact in predicting desirable actions for specific outcomes. In the latter case, we simply map the agent’s reactions to the status of the in vivo cells and environmental factors while recording the outcome of these actions. Over time, we will approximate probability distributions of the action space and the actual outcomes of these actions, facilitating analysis over the experimental data [pearlcauses] (See Appendix 6.4).
@@ 6.1.1 Induction of Heart Defects
  • RNA Interference (RNAi): Using UAS-RNAi lines targeting essential cardiac genes (e.g., hand, dMEF2), we knock down gene expression specifically in the heart, leading to structural or functional defects [Lesch2010].
@@ 6.1.2 Isolation and Maintenance of the Adult Heart Organ
The isolated heart organ is transferred to a specialized perfusion chamber designed to maintain tissue viability. The chamber provides continuous flow of oxygenated artificial hemolymph at 25°C. The heart is pinned gently to a Sylgard-coated dish to stabilize it without causing additional damage [Wasserthal2007].
A custom-designed three-dimensional (3D) microelectrode array (MEA) is fabricated to interface directly with the ex vivo heart organ. The MEA consists of microelectrodes arranged to match the geometry of the Drosophila heart, enabling localized application of electrical stimuli and recording of bioelectric signals [scalableElectro].
@@ 6.2.1 Current Understanding of Bioelectric Signals
Bioelectric signals are fundamental to the regulation of cellular and developmental processes. They are generated by the movement of ions across cellular membranes, creating voltage gradients ($V_{\text{mem}}$) that influence cell behaviour. Recent research has begun to map the bioelectric patterns associated with various cellular states and developmental processes, revealing that these patterns play a crucial role in tissue formation, regeneration, and organogenesis [Levin2009, Levin2014wh, adams2013endogenous].
Just as the genome has been sequenced to understand genetic contributions to biology, efforts have been made to map bioelectric patterns in organisms. Techniques such as voltage-sensitive dyes and fluorescent reporters have allowed visualization of membrane potential distributions in living tissues. Studies have shown that specific bioelectric patterns correlate with particular developmental outcomes. For example, the formation of the anterior-posterior axis in Xenopus laevis embryos is regulated by bioelectric cues [adams2013endogenous, Pai2012].
However, unlike the genome, the bioelectric "map" is dynamic and context-dependent. The spatial and temporal variability of bioelectric signals across different cell types and developmental stages makes comprehensive mapping challenging. Current maps are incomplete and often organism-specific, limiting the generalization of findings [McCaig2005, Sundelacruz2009].
Cells interpret bioelectric signals through voltage-gated ion channels, transporters, and voltage-sensitive signalling pathways. Changes in membrane potential can influence gene expression, cell proliferation, differentiation, migration, and apoptosis. For instance, depolarization of membrane potential has been shown to promote stem cell proliferation, while hyperpolarization can induce differentiation [Levin2009, Levin2014wh, Blackiston2011].
Despite these insights, the precise mechanisms by which cells transduce bioelectric signals into specific biochemical responses remain poorly understood. The complexity arises from the interplay between electrical, chemical, and mechanical signals within the cellular microenvironment. Additionally, the same bioelectric signal can elicit different responses depending on the cell type and its developmental context [Levin2014wh].
@@ 6.3 Introduction to Neuromorphic Computing
Neuromorphic computing is an emerging field that seeks to emulate the architecture and functioning of the biological nervous system in hardware and software systems [Indiveri2011]. These systems typically utilize spiking neural networks (SNNs) and other brain-inspired architectures to achieve highly parallel and energy-efficient processing. Unlike traditional von Neumann architectures, neuromorphic systems integrate memory and computation, potentially offering advantages in real-time data processing and low-latency responses [Benjamin2014].
@@ 6.3.1 Analogies Between Neuromorphic Systems and Cellular Bioelectric signalling
Biological cells communicate through bioelectric signals, primarily mediated by the flow of ions across cell membranes, which generate voltage gradients ($V_{\text{mem}}$) essential for various cellular functions [Levin2014wh]. Similarly, neuromorphic systems employ spiking neurons that communicate via discrete electrical pulses, akin to action potentials in biological neurons [Maass1997]. This parallel suggests a potential framework where neuromorphic hardware could model and interact with cellular bioelectric behaviours, enabling more biologically plausible simulations and controls within the DRL framework.
While the application of neuromorphic computing to biological systems is still in its nascent stages, preliminary studies indicate promising directions. For instance, [Roy2019] explored the use of SNNs to model neural dynamics in biological tissues, achieving real-time simulations of neuronal networks.

### v2

@@ 1.2 Synthetic Biology and Multicellularity
Pezzulo and Levin [34] advocate for top-down approaches in understanding and controlling complex biological systems, contrasting with traditional bottom-up, molecular-centric frameworks. Their study emphasizes the benefits of higher-level organizational strategies, akin to those in physics and engineering, for regulating complex processes such as pattern formation and regeneration. The methodology involves integrating concepts from information theory, control theory, and computational neuroscience to develop models that focus on system-wide goal states and dynamic regulation rather than precise molecular manipulation. Examples include the use of least-action principles in physics and control-theoretic models in neuroscience to achieve system stability and desired outcomes. In developmental biology, the authors propose that targeting large-scale anatomical goals can guide cellular behavior toward desired outcomes through bioelectric signaling and pattern memory. Studies on phenomena like planaria regeneration, where bioelectric interventions led to permanent anatomical changes, support the feasibility of top-down models. The results suggest that focusing on higher-level system organization can provide more effective control over complex biological processes, paving the way for innovative approaches in regenerative medicine and synthetic bioengineering by harnessing goal-directed, system-wide controls over molecular mechanisms.
@@ 1.3 Questions
Bioelectric signals in cells are, again, electrical potentials generated by the movement of ions (such as sodium, potassium, calcium, and chloride) across the cell membrane through ion channels and pumps. Bioelectric signals are represented by the voltage gradients across cell membranes ($V_{mem}$) and play an important role in developmental biology, governing processes such as tissue formation, wound healing, and organogenesis. Recent work by Michael Levin and others has demonstrated that it is possible to manipulate these bioelectric signals, showing the capacity to direct tissue regeneration in organisms like planaria and frogs [21]. Yet, many questions remain unanswered, such as:
@@ 1.4 Research Objectives
For Objective 2, for example, a study in replicability can be conducted following [35] where an open-source framework for simulating 2D computational multiphysics problems such as electrodiffusion, electro-osmosis, galvanotaxis, voltage-gated ion channels, gene regulatory networks, and biochemical reaction networks (e.g., metabolism) has been made available through BETSE (BioElectric Tissue Simulation Engine) Github.
@@ 2.1 Current Understanding of Bioelectric Signals
Bioelectric signals are fundamental to the regulation of cellular and developmental processes. They are generated by the movement of ions across cellular membranes, creating voltage gradients ($V_{\text{mem}}$) that influence cell behaviour. Recent research has begun to map the bioelectric patterns associated with various cellular states and developmental processes, revealing that these patterns play a crucial role in tissue formation, regeneration, and organogenesis [18, 21, 1].
Just as the genome has been sequenced to understand genetic contributions to biology, efforts have been made to map bioelectric patterns in organisms. Techniques such as voltage-sensitive dyes and fluorescent reporters have allowed visualization of membrane potential distributions in living tissues. Studies have shown that specific bioelectric patterns correlate with particular developmental outcomes. For example, the formation of the anterior-posterior axis in Xenopus laevis embryos is regulated by bioelectric cues [1, 30].
However, unlike the genome, the bioelectric "map" is dynamic and context-dependent. The spatial and temporal variability of bioelectric signals across different cell types and developmental stages makes comprehensive mapping challenging. Current maps are incomplete and often organism-specific, limiting the generalization of findings [27, 44].
Cells interpret bioelectric signals through voltage-gated ion channels, transporters, and voltage-sensitive signalling pathways. Changes in membrane potential can influence gene expression, cell proliferation, differentiation, migration, and apoptosis. For instance, depolarization of membrane potential has been shown to promote stem cell proliferation, while hyperpolarization can induce differentiation [18, 21, 5].
Despite these insights, the precise mechanisms by which cells transduce bioelectric signals into specific biochemical responses remain poorly understood. The complexity arises from the interplay between electrical, chemical, and mechanical signals within the cellular microenvironment. Additionally, the same bioelectric signal can elicit different responses depending on the cell type and its developmental context [21].
@@ 2.2 Bioelectricity and its Role in Topological Control
Bioelectricity, the result of ion flow across cellular membranes, acts as a fundamental signal that coordinates cell behaviour, gene expression, and tissue organization. These voltage gradients serve as a communication medium that enables cells to “know” their relative position within an organism, guiding processes like cell proliferation, differentiation, and morphogenesis [21, 48]. Levin’s experiments with planaria have demonstrated that by manipulating these bioelectric patterns, researchers can direct tissue outcomes - regenerating a planaria’s head or a tail based solely on the bioelectric state [3]. These findings suggest that bioelectric signals function as a form of "code" that governs tissue identity, growth, and maintenance.
Additionally, bioelectric signals have been shown to play a role in broader topological manipulations. For instance, in Xenopus laevis (the African clawed frog), bioelectric manipulations have been used to induce the formation of eyes in unconventional locations, illustrating the potential of bioelectric control to not only regenerate tissues but also alter and transform them in novel ways [20]. This points to the possibility that the bioelectric "fingerprint" behind an organ or tissue could be decoded, allowing for manipulation of $V_{mem}$ at a fundamental level, potentially enabling the full recovery of damaged organs, such as in cardiac arrest victims, complete overhaul of malfunctioning tissues like cancer, and perhaps even the transformation of tissues into entirely novel structures [8].
Based on the work of Souidi et al., we take the drosophila fly as our goal system for modelling heart disease and recovery induced by the proposed DRL system [43].
@@ 2.3 Biomedical and Engineering Implications
The biomedical implications of cracking the bioelectric code are many. By understanding and controlling the high-level triggers of morphogenetic subroutines, it might seem like a giant leap in imagination to picture a moment in time when we’ll be able to induce complex regenerative outcomes such as limb and organ regeneration, repair of birth defects, and even the prevention and reprogramming of diseases like cancer via bioelectric pathways. But these processes have been made possible to some extent already in vivo, both using human cells as well as in complex organisms, such as frogs [21].
Bioelectric manipulation has shown promise in modulating the tumor environment, potentially suppressing or redirecting cancerous cell behaviour [6], while optogenetic control of voltage gradients has successfully triggered regeneration in non-regenerative species [48]. Simulations, like those done with BETSE, allow us to predict and manipulate bioelectric patterns across tissues, aiding in the exploration and improvement over the control of cellular behaviour during morphogenesis [35]. Moreover, advances in synthetic biology have begun to explore self-assembling tissues with designed bioelectric circuits, paving the way for the engineering of new tissues and organs, as well as the development of novel therapeutic strategies for regenerative medicine, such as in victims of frostbite, and synthetic life [8].
@@ 3.2 Reinforcement Learning Framework
DRL has achieved remarkable success in mastering complex games with very high dimensional state representations, as demonstrated by algorithms like AlphaGo [40], AlphaZero [41], MuZero [38], OpenAI Five for Dota 2 [29], and AlphaStar for StarCraft II [46]. These algorithms have showcased the ability of RL agents to learn optimal strategies in environments with vast state and action spaces, long time horizons, and partial observability [9].
@@ 3.2.3 Modelling Heart Topology and Bioelectric Patterns
In bioelectric experiments conducted by Michael Levin’s team, similar principles are applied to study pattern formation in regenerating tissues, such as frog embryos. Levin’s group uses voltage-sensitive dyes, such as DiBAC4(3) and CC2-DMPE, to detect membrane potential differences across cells. These dyes change their fluorescence based on the membrane potential, providing real-time spatial and temporal mapping of the bioelectric state in both normal and abnormal tissues. Through fluorescent microscopy, these bioelectric states are visualized and quantified, allowing the identification of deviations in abnormal tissues compared to healthy ones [1].
To induce a normal bioelectdrcfx ric pattern in an abnormal frog embryo, ion channels are pharmacologically modulated or optogenetics is used. Optogenetics involves light-sensitive ion channels introduced into the tissue, normally through CRISP-R, enabling precise control over membrane potential with light. By manipulating the ion flow and thus the membrane potential, researchers have been able to correct the bioelectric patterns in the abnormal tissue to match those of the healthy, normal tissue. For instance, an abnormal head pattern in a frog can be corrected by restoring the normal bioelectric gradient, guiding the cells to form a correctly structured and functional head [1].
This approach, known as bioelectric modulation, effectively uses the bioelectric pattern as a template for normal tissue development and regeneration. It demonstrates that bioelectric states serve as key regulators of tissue and organ formation, enabling control over developmental processes through targeted modulation of membrane potentials [1].
@@ 4 Causal Inference in Deep Reinforcement Learning Framework
Understanding the causal relationships between bioelectric signals and cellular behaviours could aid control and manipulation of morphogenetic processes. Integrating Judea Pearl’s do-calculus into our DRL framework could allow for more precise causal inference, helping with the agent’s decision-making processes [32, 33, 31]. This graphical representation can also function as an interpretability and explainability tool, regardless of its impact in predicting desirable actions for specific outcomes. In the latter case, we simply map the agent’s reactions to the status of the in vivo cells and environmental factors while recording the outcome of these actions. Over time, we will approximate probability distributions of the action space and the actual outcomes of these actions, facilitating analysis over the experimental data [11] (See Appendix 6.4).
@@ 6.1.1 Induction of Heart Defects
  • RNA Interference (RNAi): Using UAS-RNAi lines targeting essential cardiac genes (e.g., hand, dMEF2), we knock down gene expression specifically in the heart, leading to structural or functional defects [17].
@@ 6.1.2 Isolation and Maintenance of the Adult Heart Organ
The isolated heart organ is transferred to a specialized perfusion chamber designed to maintain tissue viability. The chamber provides continuous flow of oxygenated artificial hemolymph at 25°C. The heart is pinned gently to a Sylgard-coated dish to stabilize it without causing additional damage [47].
A custom-designed three-dimensional (3D) microelectrode array (MEA) is fabricated to interface directly with the ex vivo heart organ. The MEA consists of microelectrodes arranged to match the geometry of the Drosophila heart, enabling localized application of electrical stimuli and recording of bioelectric signals [10].
@@ 6.2.1 Current Understanding of Bioelectric Signals
Bioelectric signals are fundamental to the regulation of cellular and developmental processes. They are generated by the movement of ions across cellular membranes, creating voltage gradients ($V_{\text{mem}}$) that influence cell behaviour. Recent research has begun to map the bioelectric patterns associated with various cellular states and developmental processes, revealing that these patterns play a crucial role in tissue formation, regeneration, and organogenesis [18, 21, 1].
Just as the genome has been sequenced to understand genetic contributions to biology, efforts have been made to map bioelectric patterns in organisms. Techniques such as voltage-sensitive dyes and fluorescent reporters have allowed visualization of membrane potential distributions in living tissues. Studies have shown that specific bioelectric patterns correlate with particular developmental outcomes. For example, the formation of the anterior-posterior axis in Xenopus laevis embryos is regulated by bioelectric cues [1, 30].
However, unlike the genome, the bioelectric "map" is dynamic and context-dependent. The spatial and temporal variability of bioelectric signals across different cell types and developmental stages makes comprehensive mapping challenging. Current maps are incomplete and often organism-specific, limiting the generalization of findings [27, 44].
Cells interpret bioelectric signals through voltage-gated ion channels, transporters, and voltage-sensitive signalling pathways. Changes in membrane potential can influence gene expression, cell proliferation, differentiation, migration, and apoptosis. For instance, depolarization of membrane potential has been shown to promote stem cell proliferation, while hyperpolarization can induce differentiation [18, 21, 5].
Despite these insights, the precise mechanisms by which cells transduce bioelectric signals into specific biochemical responses remain poorly understood. The complexity arises from the interplay between electrical, chemical, and mechanical signals within the cellular microenvironment. Additionally, the same bioelectric signal can elicit different responses depending on the cell type and its developmental context [21].
@@ 6.3 Introduction to Neuromorphic Computing
Neuromorphic computing is an emerging field that seeks to emulate the architecture and functioning of the biological nervous system in hardware and software systems [15]. These systems typically utilize spiking neural networks (SNNs) and other brain-inspired architectures to achieve highly parallel and energy-efficient processing. Unlike traditional von Neumann architectures, neuromorphic systems integrate memory and computation, potentially offering advantages in real-time data processing and low-latency responses [4].
@@ 6.3.1 Analogies Between Neuromorphic Systems and Cellular Bioelectric signalling
Biological cells communicate through bioelectric signals, primarily mediated by the flow of ions across cell membranes, which generate voltage gradients ($V_{\text{mem}}$) essential for various cellular functions [21]. Similarly, neuromorphic systems employ spiking neurons that communicate via discrete electrical pulses, akin to action potentials in biological neurons [26]. This parallel suggests a potential framework where neuromorphic hardware could model and interact with cellular bioelectric behaviours, enabling more biologically plausible simulations and controls within the DRL framework.
While the application of neuromorphic computing to biological systems is still in its nascent stages, preliminary studies indicate promising directions. For instance, [37] explored the use of SNNs to model neural dynamics in biological tissues, achieving real-time simulations of neuronal networks.